### v1

# Benchmark Dataset for Pore-Scale CO2-Water Interaction

Abstract
Accurately capturing the complex interaction between CO2 and water in porous media at the pore scale is essential for various geoscience applications, including carbon capture and storage (CCS). We introduce a comprehensive dataset generated from high-fidelity numerical simulations to capture the intricate interaction between CO2 and water at the pore scale. The dataset consists of 624 2D samples, each of size $512\times 512$ with a resolution of $35\mu$m, covering 100 time steps under a constant CO2 injection rate. It includes various levels of heterogeneity, represented by different grain sizes with random variation in spacing, offering a robust testbed for developing predictive models. This dataset provides high-resolution temporal and spatial information crucial for benchmarking machine learning models.
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding authors: aa2448@hw.ac.uk and h.menke@hw.ac.uk
Background & Summary
CO2 transport through porous media plays a critical role in both natural and engineered processes, including subsurface carbon sequestration [1, 2], enhanced oil recovery [3], and groundwater management [4]. The challenge lies in accurately characterizing the movement and saturation of CO2, which is influenced by the complex interactions between fluid phases and the geological heterogeneity of the porous structure [5]. As CO2 is injected into underground formations, its movement through the pore spaces of geological materials, such as sandstone or basaltic reservoirs, dictates how efficiently it can be stored over long periods. This transport process is influenced by various factors, including capillary forces and chemical interactions between CO2, brine, and the mineral matrix.
Various approaches are utilized to understand and predict CO2 transport in porous media. Laboratory techniques, such as core flooding experiments [6], yield effective bulk properties like permeability and residual saturation. Advanced imaging methods, like X-ray micro-tomography [7], allow visualization of pore-scale phenomena but have limitations, especially for dynamic processes. Numerical simulations, including lattice Boltzmann [14], pore-network modeling [15], and direct numerical simulation [22], offer more precise estimations of the fluid properties, however at a significant computational cost.
Machine learning (ML) models are emerging as valuable tools for predicting CO2 behavior in porous media, serving as efficient surrogates for computationally expensive simulations. Recent advancements highlight ML’s potential to estimate properties, like pressure build-up and saturation levels, with impressive speed and accuracy [8, 9, 10, 11, 12, 13]. The principle of these models is to learn the relationship between inputs—such as physical properties of porous media and engineering parameters—and outputs, like spatial and temporal fluid changes. Once trained on a set of representative samples, these models can generalize to predict unseen patterns, such as new permeability fields or different injection scenarios, with considerable efficiency.
However, challenges remain in terms of having a sufficient and diverse dataset for training robust models that generalize well across various scenarios. For example, current datasets often remain constrained to relatively small scales, such as maximum mesh sizes of $256\times 256$ [16, 17, 18, 19, 20, 21], which limits the ability of these models to capture fine-grained patterns necessary for accurate predictions in complex formations. Another key limitation is that most datasets designed for machine learning models focus on predicting the final state (e.g., after the injection duration) rather than capturing intermediate states [17, 18, 19]. This limitation restricts the ability of models to capture the dynamic evolution of processes over time, which is crucial for understanding CO2 transient behaviors in real-world geological scenarios.
In this paper, we introduce a high-resolution dataset designed for benchmarking machine learning models in predicting CO2 behavior during multiphase flow in porous media. The dataset comprises 624 two-dimensional samples, each of size $512\times 512$ pixels with a spatial resolution of $35\mu$m, capturing the intricate interplay between CO2 and water over 100 equally spaced temporal snapshots under a constant CO2 injection rate. A distinctive feature of this dataset is its incorporation of varying levels of heterogeneity, represented through different grain sizes, which simulate realistic geological variability. This comprehensive dataset offers critical temporal and spatial granularity, serving as a utility for developing and benchmarking machine learning models.
Methods
Geometry Preprocessing
The geometries for the simulations are created from synthetic porous media of a single solid phase and initially the pore-space is filled with water. Initially, the original 1024 $\times$ 1024 domain was cropped into four subsamples. Each subsample was then flipped along the vertical axis, increasing the overall dataset size by a factor of eight. Each of these geometries is heterogeneous, i.e., the size of grains varies across the spatial domain (intra-sample heterogeneity). In addition, there are 5 levels of heterogeneity in the dataset (inter-sample heterogeneity) as shown in Figure 1. By exposing the ML models to a range of grain size distributions and spatial configurations, the dataset enhances the model’s ability to generalize to unseen porous media.
The inter-sample heterogeneity allows the ML model to develop robust feature extraction capabilities that are invariant to changes in grain sizes and configurations. This is crucial for ensuring that the predictions remain accurate across different geological formations. Similarly, the intra-sample heterogeneity forces the model to adapt to localized variations, which is essential for capturing fine-grained details in flow and saturation patterns. The dataset contains 624 geometries, each one is of size $512\times 512$ and the physical resolution per pixel is $35\mu$m. All samples are available in HDF5 format along with the simulations.
Figure 1: Some examples of domain geometries corresponding to different patterns of heterogeneity. The heterogeneity level increases from left to right.
Multi-phase flow at the pore-scale
Understanding CO2 injection into water-filled porous media at the pore scale is critical for designing effective carbon storage strategies, especially in tight reservoirs where pore structures are highly heterogeneous and capillary forces dominate. At this scale, the interplay between fluid properties, pore geometry, and interfacial dynamics significantly influences the distribution and transport of CO2. These micro-scale interactions can lead to complex displacement patterns including snap off, coalescence, and ganglion migration that are difficult or impossible to capture with conventional Darcy-scale constitutive functions such as saturation-dependent capillary pressure and relative permeabilities. Robust Darcy-scale models however are key to predicting CO2 migration and storage efficiency.
The two-phase flow simulations in this study were conducted using GeoChemFoam [22], an advanced open-source numerical simulator developed at the Institute of GeoEnergy Engineering at Heriot-Watt University. GeoChemFoam is based on the OpenFOAM framework and is specifically designed to investigate pore-scale processes critical to energy transition and carbon storage.
GeoChemFoam uses the algebraic Volume-of-Fluid method [23] to solve multiphase flow. The velocity $\mathbf{u}$ and the pressure $p$ solve the single-field Navier-Stokes Equations (NSE):
$\nabla\cdot\mathbf{u}$ $=0\,,$ (1)
$\rho\left(\frac{\partial u}{\partial t}+u\cdot\nabla u\right)$ $=-\nabla p+\nabla\cdot S+f_{st},$ (2)
where:
  • $\rho=\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}$ is the fluid density,
  • $\mathbf{u}$ is the velocity,
  • $S$ = $\mu\left(\nabla u+\nabla u^{T}\right)$ is the viscous stress,
  • $\mu=\alpha_{1}\mu_{1}+\alpha_{2}\mu_{2}$ is the fluid viscosity,
  • $p$ is the pressure,
  • $f_{st}$ is the surface tension force,
  • $\alpha_{i}$ is the phase volume fraction, and
  • $i=1,2$ refers to the phase index.
The surface tension force is approximated using the Continuous Surface Force (CSF) model [23]:
$f_{st}=\sigma\kappa\nabla\alpha_{1},$ (3)
where:
  • $\sigma$ is the interfacial tension, and
  • $\kappa=\nabla\cdot\left(\frac{\nabla\alpha_{1}}{|\nabla\alpha_{1}|}\right)$ is the interface curvature.
The phase indicator function $\alpha_{1}$ solves the phase transport equation:
$\frac{\partial\alpha_{1}}{\partial t}+\nabla\cdot(\alpha_{1}u)+\nabla\cdot\left(\alpha_{1}\alpha_{2}u_{r}\right)=0.$ (4)
To reduce interface smearing, an artificial compression term is introduced by replacing $u_{r}$ with a compressive velocity [23].
Figure 2: Visualization of CO2 injection in porous media initially saturated with water. The CO2 is injected from the left boundary, displacing the water phase as it migrates through the pore space.
Each geometry is a domain of 512 $\times$ 512 voxels at a resolution and depth of 35 microns. We perform a two-phase flow simulation where CO2 is injected into a fully water-filled model from the left boundary, as shown in Figure 2, at a flow rate of $1\times 10^{-8}m^{3}/s$ corresponding to a capillary number of approximately $5\times 10^{-6}$. The CO2 properties are set to be $\mu_{CO_{2}}=7.37\times 10^{-8}m^{2}/s$ and $\rho_{CO_{2}}=3.84\times 10^{2}kg/m^{3}$. The water properties are $\rho_{water}=1\times 10^{3}kg/m^{3}$ and $\mu_{water}=1\times 10^{-6}m^{2}/s$, with the interfacial tension between phases at 0.03 $N/m$, and the contact angle $\theta=45\degree$. The simulation was run until a total time of 1 s with a write interval of 0.01 s and a convergence tolerance of $1\times 10^{-8}$.
In Figure 3, we show the CO2 migration pattern, for different heterogeneities, as it displaces water at different time steps. Over time, the CO2 saturation front expands, displaying distinct channelized patterns and regions of accumulation. These patterns demonstrate the interaction between capillary forces, viscous forces, and the underlying geological features. The time-lapse progression also reveals the impact of grain size and pore structure on flow dynamics, emphasizing the importance of micro-scale processes in controlling large-scale behavior. We also show the pressure, capillary pressure, and vertical velocity fields for different geometries in Figures 4, 5, and 6, respectively.
Figure 3: CO2 (yellow) displacing water in a porous media during the simulation time. Each row shows an example of the 5 heterogeneity levels in the dataset.
Figure 4: Pressure field at different injection duration. Each row shows an example of the 5 heterogeneity levels in the dataset.
Figure 5: Capillary pressure field at different injection duration. Each row shows an example of the 5 heterogeneity levels in the dataset.
Figure 6: Vertical velocity field at different injection duration. Each row shows an example of the 5 heterogeneity levels in the dataset.
Data Records
The dataset has been made available on https://doi.org/10.5061/dryad.jm63xsjn5 and is organized into 10 folders, with each of the 5 geometries having its original version and a vertically flipped version ($2\times 5=10$). The simulation samples are provided in HDF5 format, with each file including water saturation ($\alpha_{water}$), pressure ($p$), capillary pressure ($pc$), horizontal velocity ($U_{x}$), vertical velocity ($U_{y}$), and a binary image of the physical domain (where pores are denoted by 1 and grains by 0), as detailed in Table 1 which also lists the keys required to access the data. The water saturation $\alpha_{water}$ is in the range $[0,1]$; hence, the CO2 saturation field can be computed using the relation $\alpha_{CO_{2}}=(1-\alpha_{water})\times\text{img}$, where img denotes the binary physical domain. Additionally, CSV files containing values for porosity, permeability, and relative permeability are provided, with details presented in Table 2.
File Name Key Size Description
*.hdf5 Ux $100\times 512^{2}$ x-component of flow velocity
Uy $100\times 512^{2}$ y-component of flow velocity
alpha_water $100\times 512^{2}$ water saturation field over time
img $512^{2}$ physical domain
p $100\times 512^{2}$ pressure field
pc $100\times 512^{2}$ capillary pressure field
Table 1: Overview of the dataset files, including flow velocity components, pressure fields, and physical domain representations with corresponding sizes and descriptions. Keys are provided for accessing hdf5 files.
File Name Description
poroPerm.csv Time, porosity, permeability $(m^{2})$, the characteristic pore length L, the Reynolds number Re, and the Darcy velocity $U_{D}$ at the beginning of the simulation before any CO2 is injected into the model.
relperm.csv Porosity, permeability $(m^{2})$, and the capillary number of each phase (Ca1 for water and Ca2 for CO2) at the beginning of the simulation. The saturation of water $S_{w}$, the relative permeability of water $k_{rw}$, and the relative permeability of CO2 $k_{wo}$ are shown for each output timestep.
Table 2: List of files describing porosity and relative permeability values.
Technical Validation
The GeoChemFoam solver used for flow simulation has been validated against experimental data in [24]. For accurate approximation, a convergence tolerance of $1\times 10^{-8}$ was used for all samples.
Usage Notes
The dataset is available in hdf5 format, which can be easily accessed and manipulated using Python libraries such as h5py. These files contain high-resolution data, including flow velocity components, pressure fields, and physical domain representations, making them suitable for scientific and engineering applications.
Code Availability
The input files used to simulate CO2 flow is built using GeoChemFoam [22] and is available at https://github.com/ai4netzero/generating_co2_flow. The code is written in Python 3.11.9 and the list of the requirements is shown in the readme file. GeoChemFoam can be downloaded from https://github.com/GeoChemFoam/GeoChemFoam-5.1 and has been validated against experimental data in [24].
Acknowledgments
This work is funded by the Engineering and Physical Sciences Research Council’s ECO-AI Project grant (reference number EP/Y006143/1), with additional financial support from the PETRONAS Centre of Excellence in Subsurface Engineering and Energy Transition (PACESET).
Author contributions
Conceptualization and methodology, H.P.M., J.M., A.A., A.H.E.; visualization and writing, A.A., H.P.M.; formal analysis, A.A., H.P.M., A.H.E.; funding acquisition, A.H.E., F.D., H.P.M.; supervision, A.H.E, F.D., H.P.M. All authors have read and agreed to the published version of the manuscript.
Competing interests
The authors declare no competing interests.